OpenStudy (anonymous):

What is the solution to the equation? 6 + 2s – 8s = 18 A. s = –4 B. s = –2 C. s = 2 D. s = 4 @SolomonZelman

OpenStudy (solomonzelman):

add the S-s on the left.

OpenStudy (solomonzelman):

\(\large\color{black}{ 6 \color{red}{+ 2s - 8s} = 18 }\)

OpenStudy (solomonzelman):

add the red

OpenStudy (anonymous):

the answer must be what s is

OpenStudy (anonymous):

i dont understand

OpenStudy (solomonzelman):

just follow the steps: \(\large\color{black}{ 6 + 2s - 8s = 18 }\) 1) subtract \(\large\color{black}{ + 2s-8s }\). 2) subtract \(\large\color{black}{ 6 }\) from both sides. 3) divide both sides by \(\large\color{black}{ -6 }\).

OpenStudy (solomonzelman):

do whatever you are able to do from the list of steps I gave.

OpenStudy (anonymous):

is the answer -4?

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

no it is not -4

OpenStudy (solomonzelman):

can you show me some work please?

OpenStudy (anonymous):

-2

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

What is the solution to the equation? –4q + 8 = –2q – 16 A. q = –12 B. q = –4 C. q = 4 D. q = 12

OpenStudy (anonymous):

What is the solution to the equation? negative4q + 8 = negat.2q negat. 16 A. q = neg.12 B. q = neg.4 C. q = 4 D. q = 12

OpenStudy (solomonzelman):

\(\large\color{black}{ -4q+8=-2q-16 }\) ?

OpenStudy (anonymous):

yup

OpenStudy (solomonzelman):

1) add \(\large\color{black}{ 4q }\) to both sides. 2) add \(\large\color{black}{16 }\) to both sides. 3) divide both sides by \(\large\color{black}{ 2 }\).

OpenStudy (anonymous):

-4

OpenStudy (solomonzelman):

can you please show some work here in this post, so that I can tell what you are doing?

OpenStudy (anonymous):

i dont know how this works

OpenStudy (solomonzelman):

okay, lets do step by step. Will start from step 1.

OpenStudy (anonymous):

thsnk you

OpenStudy (solomonzelman):

\(\large\color{black}{ -4q+8\color{red}{+4q}=-2q-16\color{red}{+4q} }\)

OpenStudy (solomonzelman):

tell me two things: \(\large\color{black}{ -4q+8\color{red}{+4q}~~=~? }\) \(\large\color{black}{ -2q-16\color{red}{+4q} ~~=~? }\)

OpenStudy (anonymous):

neg 4 right?

OpenStudy (solomonzelman):

-4 is what?

OpenStudy (anonymous):

the first one is 8q

OpenStudy (solomonzelman):

n it is 0q+8, which is same as just 8.

OpenStudy (solomonzelman):

just likee -4+4=0, so it -4q+4q=0

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

and the second one is?

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

q

OpenStudy (solomonzelman):

nope

OpenStudy (anonymous):

?

OpenStudy (solomonzelman):

okay, what is \(\large\color{black}{ -2q\color{red}{+4q} }\) ?

OpenStudy (anonymous):

2q

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

so then, \(\large\color{black}{ -2q-16\color{red}{+4q}=2q-16 }\)

OpenStudy (anonymous):

huh

OpenStudy (solomonzelman):

so you get: \(\large\color{black}{ -4q+8=-2q-16 }\) \(\large\color{black}{ -4q+8\color{red}{+4q}=-2q-16\color{red}{+4q} }\) \(\large\color{black}{ 8=2q-16 }\)

OpenStudy (solomonzelman):

then ad 16 to both sides.

OpenStudy (solomonzelman):

\(\large\color{black}{ 8\color{blue}{+16}=2q-16\color{blue}{+16} }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ 8\color{blue}{+16}=?}\) \(\large\color{black}{ 2q-16\color{blue}{+16}=? }\)

OpenStudy (anonymous):

24

OpenStudy (anonymous):

2q

OpenStudy (solomonzelman):

yes, and the second one is?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

\(\large\color{black}{ 24=2q }\)

OpenStudy (solomonzelman):

divide both sides by 2

OpenStudy (anonymous):

12 and 1

OpenStudy (solomonzelman):

yes: \(\large\color{black}{ 12=q }\)

OpenStudy (anonymous):

thank you sooooooooo much @SolomonZelman

OpenStudy (solomonzelman):

yw!